An arithmetic series has 20 terms. The first is -50 and the last term is 83, find the sum of the series
Answer by checkley77(12844)

SUM=N/2(A+L) WHERE N=NUMBER OF TERMS, A=THE FIRST TERM, L=THE LAST TERM.
S=20/2(-50+83)
S=10*33
S=330 ANS.